OpenStudy (anonymous):

What is the solution of the equation 4x + 17 = 45?

OpenStudy (anonymous):

First of all: Subtract 17 from both the sides, and tell me what did you get??

OpenStudy (anonymous):

4x+17=45 -17 -17 4x 28 4 4 =

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Ohh ok.. i get it